OpenStudy (john_gardner):

If a Rubik’s Cube has a volume of 384 cubic centimeters, how long is one side of the cube? (Recall that the volume of a cube is calculated byl^3, where l is the length of one side.) A. 4^3√6 centimeters B.4^3√24 centimeters C. 16^3√6 centimeters D. 16^3√24 centimeters

OpenStudy (john_gardner):

help please -john

OpenStudy (supersmart1001):

Let l be side of the cube.. => its volume = l^3 here, 384 = l^3 => l = cube root of 384

OpenStudy (supersmart1001):

let x be the side vol =x*x*x x^3 = 384 x=

OpenStudy (supersmart1001):

x= ______

OpenStudy (supersmart1001):

@john_Gardner

OpenStudy (john_gardner):

Yikes too confusing :/

OpenStudy (john_gardner):

@supersmart1001 i find the cube root of 384 then i'm guessing

OpenStudy (supersmart1001):

good guess thats correct

OpenStudy (john_gardner):

I got A is that right?

OpenStudy (supersmart1001):

i got that